OpenStudy (anonymous):

f(x)= 3x- 7 and g(x)=-2x-6 find (f o g)(4)

OpenStudy (anonymous):

=(fog)(4) =f(g(4)) *replace x with 4 =f(-2(4) -6) =f(-14) *replace x with -14 =3(-14)- 7 =-42-7 = -49 so.. the answer : (f o g) (4) = -49 i think this is the solution.. :)